• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Seonho Choi

    Log entry time 23:57:50 on December 04, 2006

    Entry number 190048

    keyword=Swing Shift Summary (Mon. Dec. 4, 2006)

    16:00 Recovery in progress from ESR crash
    16:30 Called for controlled access to reset magnet power supplies that are
    gone during ESR crash
    17:10 Magnets are back. Ready for the experiment.
    17:20 Target moved to He3 position. Asked 30uA
    17:40 MCC called. They have some problems with BLM and working on it.
    17:45 Beam back. Start to take data
    Run 1311
    Run 1312 DT is about 20%
    Run 1313
    20:06 MCC called. They want to do PZT scan. Stopped run 1313
    While waiting for MCC, we changed kinematics
    L: 3.051 GeV, 19.399 deg
    R: 0.5193 GeV, 69.590 deg
    20:32 MCC finished PZT scan. Restoring 30uA beam to the hall.
    20:32 Adjusting Cerenkov HV
    Run 1314 junk
    Run 1315 Good run.
    21:10 MCC called. The beam will be stopped until 21:30. They have to
    replace crate power supply, etc.
    21:30 Beam restored to 30uA
    Run 1316 Another good run
    Run 1317 Still more good runs